# Ivy-Fake: A Unified Explainable Framework and Benchmark for Image and Video AIGC Detection

### v1

@@ Abstract
The rapid advancement of Artificial Intelligence Generated Content (AIGC) in visual domains has resulted in highly realistic synthetic images and videos, driven by sophisticated generative frameworks such as diffusion-based architectures. While these breakthroughs open substantial opportunities, they simultaneously raise critical concerns about content authenticity and integrity. Many current AIGC detection methods operate as black-box binary classifiers, which offer limited interpretability, and no approach supports detecting both images and videos in a unified framework. This dual limitation compromises model transparency, reduces trustworthiness, and hinders practical deployment. To address these challenges, we introduce Ivy-Fake , a novel, unified, and large-scale dataset specifically designed for explainable multimodal AIGC detection. Unlike prior benchmarks, which suffer from fragmented modality coverage and sparse annotations, Ivy-Fake contains over 150,000 richly annotated training samples (images and videos) and 18,700 evaluation examples, each accompanied by detailed natural-language reasoning beyond simple binary labels. Building on this, we propose Ivy Explainable Detector (Ivy-xDetector), a unified AIGC detection and explainable architecture that jointly performs explainable detection for both image and video content. Our unified vision-language model achieves state-of-the-art performance across multiple image and video detection benchmarks, highlighting the significant advancements enabled by our dataset and modeling framework. Our data is publicly available at https://huggingface.co/datasets/AI-Safeguard/Ivy-Fake.
@@ 1 Introduction
Artificial Intelligence Generated Content (AIGC) has advanced rapidly, particularly in visual modalities such as images and videos. These developments have been driven by powerful generative models, notably diffusion-based architectures like DALL-E (Betker et al., 2023), Imagen (Saharia et al., 2022), and Stable Diffusion (Rombach et al., 2022a), which have redefined state-of-the-art performance in text-to-image synthesis. More recently, models such as OpenAI’s SORA (Brooks et al., 2024) have extended AIGC into video generation, producing visually coherent, minute-long videos conditioned on text prompts with remarkable fidelity and semantic consistency.
@@
While these breakthroughs unlock transformative applications across digital media, advertising, and entertainment, they simultaneously exacerbate challenges related to content authenticity. The growing realism of synthetic content increasingly blurs the boundary between genuine and fabricated media, raising critical concerns regarding misinformation, content provenance, and public trust.
To date, most AIGC detection methods have framed the problem as a binary classification task: determining whether content is authentic or AI-generated. However, these approaches typically offer limited interpretability, which undermines transparency and user confidence. Many rely on low-level statistical cues or implicit features without providing insight into which regions of an image or video contribute to the detection outcome. Furthermore, current detection datasets exhibit constraints in terms of generator diversity, modality coverage, and annotation quality, hindering rigorous evaluation of detection models under varied and complex real-world scenarios.
In response to these limitations, Multimodal Large Language Models (MLLMs) have emerged as promising candidates for AIGC detection. By integrating visual and linguistic reasoning, MLLMs offer unique advantages, enabling not only authenticity assessment but also natural language explanations of detected anomalies. This capacity for explainability enhances the interpretability and credibility of detection results, positioning MLLMs as valuable tools for mitigating risks associated with synthetic media proliferation.
@@
Nevertheless, existing benchmarks remain inadequate for evaluating explainable AIGC detection. Datasets such as AIGCDetectionBenchmark (Zhong et al., 2023) and GenVideo (Chen et al., 2024a) provide only binary labels, while more recent resources like LOKI (Ye et al., 2025) attempt to incorporate fine-grained anomaly annotations across modalities but remain limited in scale and diversity. Similarly, FakeBench (Li et al., 2024c) emphasizes explainable fake image detection, yet omits video content, and FakeClue (Wen et al., 2025) offers extensive image-level annotations but lacks integrated video data. This fragmentation results in significant gaps in benchmarking and hinders unified progress.
@@
To address these challenges, we propose Ivy-Fake , a comprehensive benchmark designed to evaluate explainable multimodal AIGC detection. Ivy-Fake  offers: 1) Diverse Multimodal Data, a large-scale dataset comprising 94,781 annotated images and 54,967 annotated videos for training, along with 18700 samples per modality for evaluation. 2) Explainable Annotations, rich annotations that extend beyond binary labels to include detailed reasoning, enabling nuanced evaluation of models’ interpretability and explanatory capabilities. Based on our proprietary dataset, we introduce Ivy-xDetector, a model for detecting AI-generated images and videos and explaining the associated artifacts. As demonstrated by Figure 1, unlike other MLLMs, Ivy-xDetector  excels at spotting generative artifacts, spatial in images and both spatial and temporal in videos. By integrating multiple spatial and temporal feature extractors, it detects image-level artifacts and video-level temporal inconsistencies with superior accuracy. Our main contributions are summarized as follows:
  • Unified Vision-Language Detector: We introduce the first unified vision–language detector that provides detailed, explainable AIGC detection for both images and videos, achieving state-of-the-art performance across multiple detection and explainability benchmarks.
@@
  • Unified Multimodal Detection & Explainability Dataset: We present the first large-scale benchmark for explainable AIGC detection across images and videos, featuring 5.4 million labeled samples plus 150 000 instances enriched with fine-grained visual annotations and textual justifications to enable transparent, multimodal evaluation and future research.
2 Related Work
2.1 Synthetic Content Detection
Due to growing concerns about the misuse of synthetic data (Deng et al., 2024), research on AI-generated content (AIGC) detection has expanded rapidly in recent years. Most existing models for AI-generated images and videos formulate the task as binary classification, simply predicting whether the content is "real" or "fake." Representative examples include CNN-based AIGVDet (Bai et al., 2024), CNNSpot (Wang et al., 2020) and Transformer-based models such as DIRE (Wang et al., 2023) and AIDE (Yan et al., 2025). Meanwhile, several works have explored the application of multimodal large language models (MLLMs) (Zheng et al., 2024; Guo et al., 2025b; Zheng et al., 2023a) to AIGC detection, including FakeBench (Li et al., 2024c), LoKI (Ye et al., 2025), Synartifact (Cao et al., 2024) and Bi-LORA (Keita et al., 2025). However, these approaches largely overlook the importance of interpretability in AIGC detection.
Some efforts attempt to introduce interpretability by leveraging spatial annotations (Dong et al., 2022) or frequency-domain artifact analysis (Zhang et al., 2023). Nevertheless, the resulting explanations are often difficult for humans to comprehend, as they lack clarity in natural language. This limitation is particularly evident in the video domain, where AI-generated content frequently exhibits obvious flaws, e.g., incoherent frame transitions and object inconsistency, that are easily noticed and reasoned about by humans (Deng et al., 2024). FakeClue (Wen et al., 2025) introduces the use of vision-language models (VLMs) to provide interpretability for image-level detection, but it does not offer a unified framework that integrates both images and videos.
2.2 Datasets for Synthetic Content Detection.
Early datasets for synthetic content detection, such as CNNSpot (Wang et al., 2020), primarily collected fake images generated by GAN-based models (Goodfellow et al., 2014; Zhu et al., 2017; Brock et al., 2018). However, with the advent of more advanced generative architectures like diffusion models (Ho et al., 2020; Dhariwal, Nichol, 2021; Rombach et al., 2022b; Hertz et al., 2022; Nichol et al., 2021) and their variants, the authenticity of generated content has significantly increased, making it more challenging for detection models to discern. This has spurred the development of newer datasets, including ArtiFact (Cao et al., 2024), GenImage (Zhu et al., 2023b), and WildFake (Hong et al., 2025). GenImage (Zhu et al., 2023b) comprises images from the 1000 ImageNet (Russakovsky et al., 2015) categories, generated by eight state-of-the-art generators such as Stable Diffusion (Rombach et al., 2022a) and Midjourney. Nevertheless, these datasets predominantly focus on image-based content. More recently, datasets emphasizing interpretability have also been introduced. FakeClue (Wen et al., 2025) contains a large amount of image data with explainability annotations but lacks video data. LOKI (Ye et al., 2025) offers data across 26 different categories and includes 18,000 distinct questions; however, its volume of multimodal data is relatively small and primarily suited for evaluation rather than comprehensive model training. Therefore, a critical gap exists for a unified benchmark encompassing both image and video modalities to rigorously evaluate the performance of contemporary AIGC detectors.
@@ 3 Dataset
Figure 2: Overview of the proposed unified and explainable IVY-FAKE dataset. Input images or videos from diverse domains are processed alongside specific prompts by an MLLM. The model leverages temporal and spatial analysis to produce structured, explainable annotations.
The Ivy-Fake  dataset is a comprehensive dataset specifically designed for explainable multimodal AIGC detection. It contains a substantial scale of data, comprising 94,781 images and 54,967 videos in the training set, alongside 8,731 images and 9,956 videos allocated for testing. To ensure extensive coverage of potential AIGC scenarios, the dataset features significant content diversity, encompassing numerous categories such as animals, objects, human portraits, scenes, documents, satellite imagery, and DeepFake media, as shown in Figure 2. Moreover, we prioritize source diversity by including synthetic data generated through various cutting-edge architectures, i.e., GANs, Diffusion models, and Transformer-based generators, paired with authentic content from real-world contexts. Additionally, to maintain dataset relevance in the face of rapidly evolving generative technologies, we supplemented publicly available data by continuously collecting the latest AIGC images and videos from online sources. This strategy ensures that Ivy-Fake  remains current, comprehensive, and reflective of contemporary generation techniques. Next, we will detail how the Ivy-Fake  dataset is created.
@@ 3.1 Data Collection
3.1.1 Video Dataset Construction
Public Benchmark Sources A total of approximately 110,000 video clips were collected from diverse sources to ensure comprehensive coverage of contemporary generative models. A significant portion was sourced from established public datasets, notably GenVideo (Chen et al., 2024a) and LOKI (Ye et al., 2025), which offer large collections of AI-generated videos alongside authentic counterparts. These datasets include outputs from a variety of advanced video generation frameworks, ensuring broad model diversity. Specifically, our dataset includes samples generated by state-of-the-art models such as SORA (soravideo2024), Stable Video Diffusion (SVD), Runway, ModelScope, Latte, VideoCrafter, and OpenSora.
Web-Crawled Sources To supplement public datasets and enhance real-world relevance, we crawled additional video content from publicly accessible platforms. This included authentic and suspected AI-generated videos from popular sources such as YouTube and various social media platforms. This web-crawled content contributes valuable diversity and reflects a broader spectrum of naturally occurring AI-generated media.
3.1.2 Image Dataset Construction
Public Benchmark Sources Approximately 110,000 images were collected following a similar methodology to the video dataset. Publicly available datasets, including FakeClue (Wen et al., 2025) and WildFake (Hong et al., 2025), were utilized to gather a wide range of AI-generated images. These benchmarks include content produced by mainstream generative models, notably GANs (Goodfellow et al., 2014) and Diffusion models, thus ensuring the inclusion of diverse generative paradigms.
Web-Crawled Sources To further enrich the dataset and capture authentic and synthetic images prevalent in real-world scenarios, additional samples were sourced from various public websites. These included social media platforms, image-sharing sites, and other third-party repositories that host both genuine and AI-generated images.
3.2 Sampling Strategy and Dataset Balance
@@ 3.3 Data Annotation
To generate explainable annotations, we employed the multimodal large language model Gemini 2.5 Pro (Team et al., 2023), leveraging a knowledge distillation process to produce structured, interpretable outputs. Each of the collected images and videos is classified as either authentic or AI-generated. The specific system prompt guiding this task is provided in Appendix. To ensure consistent and structured output, we adapted the distillation template from DeepSeek-R1-Zero (Guo et al., 2025a), which requires the model to first articulate a reasoning process before issuing a final determination, as outlined in Figure 2. This structured approach was essential to mitigate content-related biases, such as enforcing reflective reasoning or imposing predefined problem-solving patterns, thereby allowing the model’s natural reasoning process to emerge during reinforcement learning (RL). Furthermore, this format facilitated the identification and exclusion of non-compliant distillations, ensuring high data quality and consistency for subsequent model training.As noted in FakeVLM (Wen et al., 2025), directly predicting content authenticity using vision-language models exhibits performance limitations. FakeVLM addresses this by adopting a more complex output structure instead of a simple “Real”/“Fake” binary classification. In contrast, our approach introduces a structured response format using <think></think> and <conclusion></conclusion> tags. This template guides the model to first articulate its reasoning process before arriving at a final decision, which has been empirically shown to improve both detection accuracy and interpretability.
To enhance annotation accuracy and guide the generation of relevant explanations, we incorporated category-level prior knowledge during annotation. Specifically, Gemini was provided with the ground truth authenticity label (i.e., "real" or "fake") and instructed to explain the rationale behind this classification. The user prompt was standardized as follows:
This {file_type} is {label}. Explain the reason.
In this template, {file_type} indicates the modality of the input—either "image" or "video"—and {label} represents the ground-truth label, assigned as either "real" or "fake". The distilled explanations were further categorized along two major dimensions to facilitate structured analysis: Spatial Features, which comprises eight sub-dimensions and captures artifacts and inconsistencies observable within individual frames or static images. Temporal Features, which includes four sub-dimensions and describes anomalies associated with motion, temporal coherence, and cross-frame consistency. Since still images inherently lack temporal attributes, this category is exclusively applicable to video annotations. These categories were informed by established taxonomies of visual artifacts in generative content, as detailed in prior research (Deng et al., 2024).

### v2

@@ Abstract
The rapid development of Artificial Intelligence Generated Content (AIGC) techniques has enabled the creation of high-quality synthetic content, but it also raises significant security concerns. Current detection methods face two major limitations: (1) the lack of multidimensional explainable datasets for generated images and videos. Existing open-source datasets (e.g., WildFake, GenVideo) rely on oversimplified binary annotations, which restrict the explainability and trustworthiness of trained detectors. (2) Prior MLLM-based forgery detectors (e.g., FakeVLM) exhibit insufficiently fine-grained interpretability in their step-by-step reasoning, which hinders reliable localization and explanation. To address these challenges, we introduce Ivy-Fake, the first large-scale multimodal benchmark for explainable AIGC detection. It consists of over 106K richly annotated training samples (images and videos) and 5,000 manually verified evaluation examples, sourced from multiple generative models and real-world datasets through a carefully designed pipeline to ensure both diversity and quality. Furthermore, we propose Ivy-xDetector, a reinforcement learning model based on Group Relative Policy Optimization (GRPO), capable of producing explainable reasoning chains and achieving robust performance across multiple synthetic content detection benchmarks. Extensive experiments demonstrate the superiority of our dataset and confirm the effectiveness of our approach. Notably, our method improves performance on GenImage from 86.88% to 96.32%, surpassing prior state-of-the-art methods by a clear margin.
00footnotetext: $\clubsuit$ Equal Contributions. 🖂Corresponding author
@@ 1 Introduction
The rapid development of diffusion-based models has triggered an exponential growth in Artificial Intelligence Generated Content (AIGC), like Sora [sora_videoworldsimulators2024], DALL-E [dalle_3_betker2023improving], Imagen [imagegen_saharia2022photorealistic], and Stable Diffusion [rombach2022high], which have redefined state-of-the-art performance in text-to-image synthesis. However, these advances also raise significant security concerns, including the misuse of DeepFake [Qureshi2024DeepfakeFA], document tampering [Qureshi2024DeepfakeFA], and dataset poisoning [10.5555/3294996.3295110]. The increasing realism of synthetic content blurs the boundary between genuine and fabricated media, posing critical challenges for misinformation control, content provenance verification, and the preservation of public trust.
@@
However, most existing approaches primarily focus on binary authenticity classification [demamba, wildfake_2025, aigvdet_bai2024ai, aide_yan2025a], which limits the human interpretability of model predictions [loki_ye2024loki]. Existing benchmarks also remain inadequate for evaluating explainable AIGC detection. Datasets such as AIGCDetectionBenchmark [aigcdetectionbenchmark_zhong2023patchcraft] and GenVideo [demamba] provide only binary labels, while more recent resources like LOKI [loki_ye2024loki] attempt to incorporate fine-grained anomaly annotations across modalities but are still constrained in scale and diversity. On the other hand, current methods such as AIDE [aide_yan2025a] and Demamba [demamba] are limited to a single modality—either image or video—without exploring unified detection across both domains. Similarly, existing datasets exhibit the same fragmentation: FakeBench [liu2024mmfakebench] emphasizes explainable fake image detection but omits video content, whereas FakeClue [fakeclue_wen2025spot] provides extensive image-level annotations yet lacks integrated video data. This fragmentation leads to substantial gaps in benchmarking and hinders unified advancement in explainable multimodal AIGC detection.
@@
To address these challenges, we propose Ivy-Fake, a comprehensive dataset designed to evaluate explainable multimodal AIGC detection. Ivy-Fake offers: 1) Diverse Multimodal Data, a large-scale dataset comprising 61,107 annotated images and 45,272 annotated videos for training, along with 5,000 samples for evaluation. 2) Explainable Annotations, rich annotations that extend beyond binary labels to include detailed reasoning, enabling nuanced evaluation of models’ interpretability and explanatory capabilities.
In addition to providing Ivy-Fake, we introduce Ivy-xDetector, a model for detecting AI-generated images and videos and explaining the associated artifacts. As demonstrated by Figure 1, unlike other MLLMs, Ivy-xDetector excels at spotting generative artifacts, spatial in images and both spatial and temporal in videos. By integrating multiple spatial and temporal feature extractors, it detects image-level artifacts and video-level temporal inconsistencies with superior accuracy. Through the incorporation of explainable annotations and carefully curated domain-diverse data, our approach achieves state-of-the-art performance using only a fraction of the data required by existing detectors. While prior image/video classifiers [demamba, aide_yan2025a, aigcdetectionbenchmark_zhong2023patchcraft] often rely on millions (M-level) of samples, our method attains superior performance across both image and video AIGC detection tasks with fewer than 200K samples.
Our main contributions are summarized as follows:
@@
  • We introduce Ivy-Fake, the first large-scale dataset for explainable AIGC detection across both images and videos, comprising 106,379 training samples and 5,000 manually verified test instances. Each entry is enriched with fine-grained visual annotations and textual reasoning to support transparent multimodal evaluation and future research.
  • We propose Ivy-xDetector, a reinforcement learning–based model by Reinforcement Learning with Verifiable Rewards (RLVR) built upon GRPO, which achieves superior performance on multiple image and video detection benchmarks using fewer than 200K samples—significantly outperforming existing methods that rely on millions of data points.
  • We conduct extensive experiments to validate the effectiveness of both the dataset and the model, demonstrating their substantial impact on improving multimodal large models’ capability in synthetic content detection.
@@ 2.1 Data Collection
The Ivy-Fake dataset is curated from GenVideo [demamba], LOKI [loki_ye2024loki], FakeClue [fakeclue_wen2025spot], WildFake [wildfake_2025], Kinetics-400 [kay2017kineticshumanactionvideo], and MMFakeBench [liu2024mmfakebench]. Since Ivy-Fake is substantially larger than other datasets, we primarily detail our data processing procedures here.
Image/Video Corruption Filtering.
The integrity of images and videos is critical for MLLM training. We observed that the official GenVideo [demamba] repository contained a large number of unreadable videos. To ensure data quality, we employed the official Qwen2.5-VL [bai2025qwen2] I/O library to filter out corrupted files, discarding images and videos that failed to load.
Image/Video Resolution.
To maintain consistency, we only retained images whose resolutions fell within [3,136, 12,845,056] pixels, and videos whose per-frame resolutions were within [100,352, 602,112] pixels, with the additional constraint that total pixels did not exceed 19,267,968. Samples outside these ranges were discarded.
Table 1: Task examples of Ivy-Fake.
Features Sub Dimension Example
Spatial Impractical Which visual clue reveals an impossible spatial or lighting configuration in the scene?
Luminosity (A) Light direction opposite source. (B) Face without lighting. (C) Dark background, bright face. (D) Bright background, dark face.
Localized Which clue shows local blur or focus inconsistency?
Blur (A) Clear face but blurred hand. (B) Sharp background, fuzzy hair. (C) Glasses edges blurred, eyes clear. (D) Hand blurred while holding clear object.
Illegible Which part of the document text is tampered or altered?
Letters (A) Text blurred or half missing. (B) Letters inconsistent in size. (C) Words overlap or melt together. (D) Signs contain random symbols.
Omitted Which part of the person looks incomplete or missing?
Components (A) Missing one finger. (B) Two fingers merged. (C) Missing eyebrows. (D) Three pairs of glasses.
Spatial Where does the spatial relation look unnatural or physically impossible?
Relationships (A) Arm passes through body. (B) Legs overlap unnaturally. (C) Head detached from neck. (D) Hand holds object from wrong side.
Distorted Which part appears deformed or stretched in shape?
Components (A) Finger bent at odd angle. (B) Eye stretched sideways. (C) Mouth warped or uneven. (D) Hand larger than face.
Chromatic Which clue shows unnatural or inconsistent colors in the image?
Irregularity (A) Face partly blue under warm light. (B) Green reflection on red shirt. (C) Sky color bleeds into hair. (D) Background color unevenly saturated.
Abnormal What visual area shows abnormal or inconsistent texture?
Texture (A) Skin looks like plastic. (B) Hair pattern repeats unnaturally. (C) Cloth surface too smooth or glossy. (D) Wall texture mixed with skin.
Temporal Luminance Which moment shows sudden or inconsistent brightness between frames?
Discrepancy (A) Light flickers without reason. (B) Face turns dark while background stays bright. (C) Shadow jumps between frames. (D) Object brightness changes abruptly.
Awkward When does the facial movement look unnatural or discontinuous across frames?
Facial Expression (A) Smile appears and disappears suddenly. (B) Eyes blink in different directions. (C) Mouth freezes mid-motion. (D) Expression changes without emotion cue.
Duplicated Where does an object or body part appear twice across frames?
Component (A) Two faces shown in one frame. (B) Hand shadow moves separately. (C) Extra arm appears then fades. (D) Same object duplicated in next frame.
Non-Spatial When does the action sequence look illogical or out of order?
Relationships (A) Person closes door before touching it. (B) Cup moves without contact. (C) Smile appears after laughter. (D) Object vanishes then reappears suddenly.
Near-Duplicate Image Filtering.
Selecting challenging and diverse samples is crucial for model robustness. We extracted visual features from images and video frames using CLIP [Radford2021LearningTV] and computed intra-class similarity scores. Images with excessively high similarity were removed, while videos were left unprocessed. To comprehensively activate and enhance the model’s reasoning capabilities across the task, we propose a high-quality data construction pipeline that integrates explainable detection of AI-generated content.
Public Dataset Sources
The Ivy-Fake training set comprises approximately 106K image and video samples collected during the cold-start phase from diverse sources to ensure broad coverage of contemporary generative models. A significant portion was sourced from established public datasets, including GenVideo [demamba], LOKI [loki_ye2024loki], FakeClue [fakeclue_wen2025spot], WildFake [wildfake_2025], Kinetics-400 [kay2017kineticshumanactionvideo], MMFakeBench [liu2024mmfakebench], GenImage [genimage_2023], DiffusionForensics [dire_wang2023dire]. The cold-start phase aimed to enhance the granularity of interpretable content. To further improve the detector’s accuracy, an additional 87,500 image/video samples were stratified and sampled from multiple datasets for RLVR training, including MSRVTT [xu2016msrvtt] (9,000 videos), GenImage [genimage_2023] (32,000 images per category), CnnSpot [cnnspot_wang2020cnn] (19,000 samples), DigiFakeAV [liu2025faceswappingdiffusionbaseddigital] (2,000 samples), GenVideo [demamba] (Fake subset) (9,000 samples), DiffusionForensics [dire_wang2023dire] (4,500 samples), WildFake [wildfake_2025] (12,000 samples), and synthetic data generated by Qwen-Image [wu2025qwenimagetechnicalreport] (634 samples). In total, the dataset used for training comprises 194,513 samples.
Generated Sources
To further enrich the dataset and better capture synthetic images prevalent in real-world scenarios, additional sadmples were generated using a diverse set of state-of-the-art image generation models. These models span multiple architectures and training paradigms, ensuring broad coverage of visual styles and generative patterns for robust evaluation. Specifically, extra sources includes samples generated by FLUX.1 [labs2025flux1kontextflowmatching], Lora Flux [labs2025flux1kontextflowmatching] and Qwen-Image [wu2025qwenimagetechnicalreport].
@@ 2.2 Data Annotation
Figure 2: Token Length Distributions and Multi-Dimensional Coverage Across Datasets. Left: Distribution of token lengths across datasets; Right: Coverage of multiple dimensions in explainability datasets, extracted using Qwen3-32B [qwen3]. The Prompt can be seen in appendix.
To generate explainable annotations, we employed the multimodal large language model Gemini 2.5 Pro [genimi_team2023gemini], leveraging a knowledge distillation process to produce structured, interpretable outputs. To ensure consistent and structured output, we adapted the distillation template from DeepSeek-R1-Zero [deepseekr1_guo2025deepseek], which requires the model to first articulate a reasoning process before issuing a final determination.
To enhance annotation accuracy and guide the generation of relevant explanations, we incorporated category-level prior knowledge during annotation. Specifically, Gemini was provided with the ground truth authenticity label (i.e., "real" or "fake") and instructed to explain the rationale behind this classification. The user prompt was standardized as follows: This {file_type} is {label}. Explain the reason. To further ensure the quality and consistency of explainable annotations, we conducted a comprehensive human alignment and verification process, as detailed in Appendix 9.
As illustrated in Table 1, file_type indicates the modality of the input—either "image" or "video"—and {label} represents the ground-truth label, assigned as either "real" or "fake". The distilled explanations were further categorized along two major dimensions to facilitate structured analysis: Spatial Features, which comprises eight sub-dimensions and captures artifacts and inconsistencies observable within individual frames or static images. Temporal Features, which includes four sub-dimensions and describes anomalies associated with motion, temporal coherence, and cross-frame consistency. Since still images inherently lack temporal attributes, this category is exclusively applicable to video annotations. These categories were informed by established taxonomies of visual artifacts in generative content, as detailed in prior research [deng2024survey].